### v1

# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

Abstract
In this paper, we propose a coordinated routing strategy aimed at improving bus schedule adherence and enhancing travel efficiency for connected and automated vehicles (CAVs) operating within a mixed-traffic urban network. Our approach capitalizes on the existence of dedicated lanes for buses and CAVs, leveraging real-time traffic data to dynamically reroute CAVs in anticipation of congestion. By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, we enable the system to proactively adjust CAV routes when potential interference with bus operations is detected. This coordination mitigates delays affecting transit services and reduces travel time for CAVs. We evaluate the proposed strategy through simulation studies conducted in the SUMO. The results demonstrate significant improvements in both transit reliability and CAV operational performance across a range of traffic conditions.
I Introduction
Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems. Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2]. The adoption of CAVs is projected to substantially increase in the coming decades; the transition towards a fully autonomous fleet will occur gradually [3]. Transportation systems will experience prolonged periods of mixed traffic conditions where CAVs and human-driven vehicles (HVs) coexist. Understanding how to optimize the efficiency under increasing penetration rates of CAVs during this transitional period is critical [4, 5, 6]. Extensive experimental studies have indicated that heterogeneous traffic environments involving CAVs can effectively reduce total travel times [7, 8, 9]. Implementing suitable right-of-way (ROW) management strategies is essential for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transitional phase.
CAV dedicated lane (DL) has been proposed as an effective strategy to enhance the benefits that CAVs can bring to transportation systems [10]. By providing a separate driving environment, DLs enable CAVs to operate with shorter reaction times and smaller headways, thus increasing lane capacity by up to three times compared to general-purpose lanes (GPLs) [11]. In addition, DLs isolate CAVs from the mixed traffic network, offering a safer driving environment for both CAVs and HVs. This contributes to reducing vehicles’ lane changing and overtaking maneuvers, which are among the leading causes of collisions and delays in mixed traffic environments [12]. Despite the promising benefits of DLs, the high costs and limited feasibility of constructing new lanes in urban areas, due to substantial construction expenses and restricted urban space, often outweigh the advantages of CAV DLs [13]. In this context, converting an existing GPL into a CAV DL tends to be a more practical approach. However, it may lead to substantial delays for other types of vehicles using GPLs due to the reduced road capacity [14]. Recent research has shown that the overall system benefit from converting a GPL to a CAV DL may be lower than maintaining the original configuration [15].
To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL. Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [16]. While these lanes provide dedicated driving environments for transit operations, the reserved space cannot taken by other non-transit vehicle lower the traffic efficiency  [17]. Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability. In the existing literature, Luo et al. developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [18]. Chen et al. explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [19]. A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, were investigated in [11]. While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
In this paper, we propose to develop a dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses. The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency. By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs. The contributions of this work are summarized as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules. (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected. Simulation performed in SUMO demonstrates the effectiveness of the proposed approach, indicating significant improvements in traffic efficiency through coordinated use of joint DLs.
II Problem Statement
Figure 1: The road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges. Taking the road network of Manhattan as an example, as shown in Fig. 1, the node set $\mathcal{V}$ is comprised of two subsets, denoted as $\mathcal{V}\!=\!\mathcal{S}\cup\mathcal{I}$, where $\mathcal{S}$ represents the set of bus stations and $\mathcal{I}$ represents the set of road intersections. Each directed edge $(v,v^{\prime})\in\mathcal{E}$ models a road segment from node $v$ to node $v^{\prime}$, with $v,v^{\prime}\in\mathcal{V}$. Consider each edge consisting of two lanes in the same direction, i.e., joint DL and GPL, where the GPL can be used by both CAVs and HVs, while DLs are reserved only for joint use by buses and CAVs.
Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses. The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively. Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$. As illustrated in Fig. 2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable. In contrast, HVs can use only GPLs for travel purposes and route planning. With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses. Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, thus improving road capacity while avoiding disturbance caused by other HVs on GPLs.
Figure 2: Illustration of the general-purpose lane and dedicated lane.
For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented as a sequence of edges as follows:
$P_{b}=\big{\{}(v_{b,1},v_{b,2}),(v_{b,2},v_{b,3}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big{\}},$ (1)
where $v_{b,k}\!\in\!\mathcal{I}$ with $k\!=\!1,\dots,N_{b}$ denotes the $k$th road intersection along the bus route $P_{b}$, and $N_{b}$ represents the total number of intersections on the route. We refer to $v_{b,k}^{s}$ the $k$th bus stop on the edge $(v_{b,k},v_{b,k+1})$.
For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ entering the road network, let $(o_{i},d_{i})$ denote its origin-destination (OD) pair, where $o_{i},d_{i}\!\in\!\mathcal{I}$. Based on real-time traffic conditions, CAV $i$ is assigned an optimal route from its location to the destination. The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted by
$P_{o_{i},d_{i}}=\big{\{}(o_{i},v_{k}),(v_{k},v_{k+1}),\dots,(v_{n},d_{i})\big{\}},$ (2)
where $v_{k},v_{k+1},\dots,v_{n}\!\in\!\mathcal{I}$ are road intersections in the pre-planned route of the CAV. Note that, the route of each CAV is dynamically updated upon its arrival at each intersection in its route. While CAV routes may traverse both intersections and bus stations, re-optimization of the route toward the destination is triggered only when approaching an intersection. HVs are treated as passive participants in the network, following certain routes during their trips.
In this paper, we propose to develop a real-time coordinated routing approach for CAVs that fully leverages dedicated lanes, improves bus schedule adherence, and enhances CAV travel efficiency without creating new obstacles for timely bus arrivals at stops. To this end, a dynamic transportation system is modeled based on real-time traffic flow on each edge. Furthermore, CAVs are selected and rerouted to ensure on-time bus arrivals and maintain traffic efficiency for CAVs.
III Modeling of Traffic Flow
This section presents the modeling of road traffic flow in a dynamic transportation network. We start by introducing the traffic flow on joint DLs to capture the potential delays caused by CAVs. Then, the traffic flow on GPLs is modeled for effective rerouting of CAVs.
III-A Traffic Flow on Joint Dedicated Lane
The travel time for a vehicle to traverse an edge depends on the real-time traffic flow along the edge. As the number of vehicles traveling on an edge increases, traffic flow becomes heavier and travel times tend to rise due to congestion. This nonlinear relationship is well captured by the classic Bureau of Public Roads (BPR) function [20]. Let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$. According to the BPR model, the actual travel time to traverse $e$ is denoted as
$\tau_{e}=\tau_{e}^{0}\left(1+\alpha\left(\frac{f_{e}}{c_{e}}\right)^{\beta}\right),$ (3)
where $f_{e}$ represents the real-time flow on edge $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion. Both $f_{e}$ and $c_{e}$ are measured in the number of vehicles per unit time. In this paper, we assume that all vehicles, regardless of type, travel at the same speed under free-flow conditions. Additionally, $\alpha$ and $\beta$ are congestion sensitivity parameters. Since CAVs react faster and drive closer with shorter headways, they usually smooth traffic flow compared to a fully human-driven lane [18]. This results in a lower value of $\alpha$. In contrast, buses are heavier and less flexible, causing greater congestion when traffic is dense [21]. Based on these considerations, we adopt the following parameter settings in this study
$\begin{cases}\alpha\!=\!0.1,\ \beta\!=\!3,&\text{for GPL},\\ \alpha\!=\!0.2,\ \beta\!=\!5,&\text{for joint DL}.\end{cases}$ (4)
Sensors deployed in the road network enable real-time monitoring and analysis of traffic flow patterns [22]. In the road network considered, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge. For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, the traffic flow monitor window is defined as
$\mathcal{M}(t_{\text{sys}},\Delta T)=\big{[}t_{\text{sys}}\!-\!\Delta T,\;t_{\text{sys}}\!+\!\Delta T\big{]},$ (5)
where $t_{\text{sys}}$ denotes the current system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window where $\mathbb{N}_{+}$ denotes the set of positive integers. For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the transportation system at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its routing planning, which is represented as
$P_{i}(t_{\text{sys}})=\big{\{}(v_{i,1},v_{i,2}),(v_{i,2},v_{i,3}),\dots,(v_{i,n_{i}},d_{i})\big{\}},$
where $v_{i,\ell}\!\in\!\mathcal{I}$, $\ell\!=\!1,\dots,n_{i}$, denotes the $\ell$th road intersection along the planned route of CAV $i$ toward its destination. To alleviate traffic congestion on DLs that could lead to delays in bus arrivals at their stops, the cumulative flow of CAVs entering each edge on the DLs is actively monitored. Specifically, let $\tau_{i,\ell}^{0}$ represent the free-flow travel time on edge $(v_{i,\ell},v_{i,\ell+1})$, and $t_{i,\ell}$ denote the arrival time of CAV $i$ at intersection $v_{i,\ell}$. Thus, the arrival time of CAV $i$ at the next intersection $v_{i,\ell+1}$ is of the form
$t_{i,\ell+1}=t_{i,\ell}+\tau_{i,\ell}^{0}.$ (6)
This allows us to compute the accumulated number of CAVs expected to enter the edge $(v_{i,\ell+1},v_{i,\ell+2})$. To facilitate this, we define the following indicated function:
$I_{i}(t_{\text{sys}},\Delta T,v_{b,k})\!=\!\begin{cases}1,&\text{if }t_{i,\ell+1}\!\in\!\mathcal{M}(t_{\text{sys}},\Delta T),\\ &(v_{i,\ell+1},v_{i,\ell+2})\!=\!(v_{b,k},v_{b,k+1}),\\ \\ 0,&\text{otherwise}.\end{cases}$ (7)
Then, based on the number of CAVs expected to enter the edge $(v_{b,k},v_{b,k+1})$, the traffic flow on the $k$th edge in the DL of bus $b$ anticipated at time $t_{\text{sys}}$ is computed as
$\hat{f}_{b,k}(t_{\text{sys}})=\frac{\sum_{i\in\mathcal{N}^{\text{cav}}}I_{i}(t_{\text{sys}},\Delta T,v_{b,k})}{2\Delta T}.$ (8)
III-B Traffic Flow on General-Purpose Lane
Observing traffic flow on GPLs is essential to dynamically optimize CAV route planning based on real-time traffic data, thereby enhancing overall traffic efficiency. Similarly, the time window for traffic flow monitoring at the starting point of each edge on the GPLs is denoted as
$\mathcal{M}(t_{\text{sys}},\Delta\tilde{T})=\big{[}t_{\text{sys}}\!-\!\Delta\tilde{T},\;t_{\text{sys}}\!+\!\Delta\tilde{T}\big{]},$ (9)
where $\Delta\tilde{T}\!\in\!\mathbb{N}_{+}$ represents the size of the time window to observe traffic flow. With the knowledge of the route planning of CAVs at time $t_{\text{sys}}$, the anticipated traffic flow on each edge belonging to GPLs is depicted by
$\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})=\frac{\sum_{i\in\mathcal{N}^{\text{cav}}}\!I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})+\Delta N^{\text{hv}}_{(v,v^{\prime})}(t_{\text{sys}})}{2\Delta{\tilde{T}}},$ (10)
where the indicated function $I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})$ characterizes the number of CAVs anticipated to traverse the edge $(v,v^{\prime})$ within the monitoring time window. Mathematically, it is denoted as
$I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})=\begin{cases}1,&\text{if }t_{i,\ell+1}\!\in\!\mathcal{M}(t_{\text{sys}},\Delta\tilde{T}),\\ &(v_{i,\ell+1},v_{i,\ell+2})\!=\!(v,v^{\prime}),\\ \\ 0,&\text{otherwise},\end{cases}$ (11)
where $(v,v^{\prime})\!\in\!\mathcal{E}\!\setminus\!\cup_{b\in\mathcal{N}^{\text{cav}}}P_{b}$ denotes an edge located on the GPLs. Note that $\Delta{N}^{\text{hv}}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
Subsequently, according to the BPR model given in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right).$ (12)
Here, $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on edge $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow on the edge at time $t_{\text{sys}}$, computed according to (10). The congestion sensitivity parameters $\alpha$ and $\beta$ are determined by (4).
Remark 1
The traffic flow monitoring windows are defined by $\Delta{T}$ in (5) for joint DLs and by $\Delta{\tilde{T}}$ in (9) for GPLs, respectively, to capture real-time traffic conditions. A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges. With different emphasis, the sizes of the two monitoring windows are not necessarily to be identical.
IV Dynamic Rerouting Approach
In this section, we propose a coordinated rerouting approach that dynamically optimizes route assignments for CAVs based on real-time traffic data. The objective is to ensure that buses adhere to their scheduled arrival times at each stop while enhancing the overall traffic efficiency of CAVs. The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and then determining the optimal alternative routes for those CAVs.
IV-A Identification of CAVs for Rerouting
Let $\hat{t}_{b,k}$ denote the estimated arrival time of bus $b\!\in\!\mathcal{N}^{\text{b}}$ at its $k$th road intersection $v_{b,k}$. Upon arriving at its $(k\!-\!1)$th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
$P_{b}(t_{\text{sys}})=\big{\{}(v_{b,k},v_{b,k+1}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big{\}},$ (13)
excluding the intersection $v_{b,k-1}$. As such, the estimated arrival time $\hat{t}_{b,k}$ is represented as
$\hat{t}_{b,k}=t_{b,k-1}+\tau_{b,k-1}^{0},$ (14)
where $\tau_{b,k-1}^{0}$ denotes the free-flow travel time from intersection $v_{b,k-1}$ to intersection $v_{b,k}$.
Recall that the traffic flow $\hat{f}_{b,k}(t_{\text{sys}})$ on edge $(v_{b,k},v_{b,k+1})$ anticipated at time $t_{\text{sys}}$; cf. (8). Employing the BPR function in (3), the anticipated travel time on the same edge is as the following form:
$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right),$ (15)
where $c_{b,k}$ is the effective capacity of edge $(v_{b,k},v_{b,k+1})$, defined as the maximum traffic flow that the road segment can accommodate under ideal conditions before congestion affects travel times. For edges on DLs, the parameters $\alpha$ and $\beta$ are chosen following the second case in (4).
Figure 3: Illustration of the rerouting scheme at each intersection in DLs.
To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $[t_{b,k-1},\hat{t}_{b,k}]$. The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected. We make use of Fig. 3 for better illustration.
For any system time $t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]$, the set of CAVs on edge $(v_{b,k-1},v_{b,k})$ that need to be rerouted, denoted as $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, is expressed as
$\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})=\big{\{}i\!\in\!{\mathcal{N}^{\text{cav}}}:I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1\big{\}}.$ (16)
As defined (7), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of the intersection $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered. Namely,
$t_{\text{sys}}^{\prime}\in\big{\{}t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!{\tau_{b,k}^{0}}\big{\}}.$ (17)
Remark 2
In the developed rerouting approach, the traffic flow on edge $(v_{b,k},v_{b,k+1})$ is dynamically monitored, and the CAVs on the preceding edge, i.e., $(v_{b,k-1},v_{b,k})$, are controlled before bus $b$ entering $(v_{b,k},v_{b,k+1})$. This procedure enables effective congestion mitigation on each edge along the DLs. Thus, it is reasonable to employ the free-flow travel time in (14) to estimate the bus’s arrival time at $v_{b,k}$.
IV-B Route Optimization for Rerouted CAVs
To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\geq{\tau_{b,k}^{0}}$ is detected. Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to the destination $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$. Based on the latest traffic data, including estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
$\min\ \sum_{i\in\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})}\sum_{(v_{i,\ell},v_{i,\ell+1})\in{P_{i}(t_{\text{sys}}^{\prime}})}\!\!\!\!\!\!\!\hat{\tau}_{i,\ell}(t_{\text{sys}}^{\prime})$ (18)
$\ \mathrm{s.\,t.}\quad~{}P_{i}(t_{\text{sys}}^{\prime})\!\in\!\mathcal{P}_{(v_{b,k},d_{i})},~{}i\in\mathcal{R}_{b,k-1}(t_{\text{sys}^{\prime}}),$ (19)
[…]
where $\hat{\tau}_{i,\ell}(t_{\text{sys}}^{\prime})$ represents the anticipated travel time to traverse edge $(v_{i,\ell},v_{i,\ell+1})$, estimated based on traffic conditions at time $t_{\text{sys}}^{\prime}$. As defined previously, $P_{i}(t_{\text{sys}}^{\prime})$ denotes the route planning of CAV $i$ at time $t_{\text{sys}}^{\prime}$ to complete the remainder of its trip. For CAVs to be rerouted, $P_{i}(t_{\text{sys}}^{\prime})$ refers to a route from intersection $v_{b,k}$ to the destination $d_{i}$. We denote by $\mathcal{P}_{(v_{b,k},d_{i})}$ the set of all feasible routes connecting $v_{b,k}$ to $d_{i}$.
Constraints (5)-(8), (14), and (15) provide estimates of traffic flow and travel times for each edge on joint DLs, while constraints (9)-(12) offer travel time estimates for edges on GPLs. Moreover, (16) and (17) specify the triggering condition for rerouting and identify the set of CAVs selected for route replanning. After all estimated travel times across the network are updated at time $t_{\text{sys}}^{\prime}$, a prediction-aware Dijkstra algorithm [23] is executed to determine the optimal rerouting paths for all identified CAVs, denoted as $P_{i}^{*}(t_{\text{sys}}^{\prime})$, where $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$. In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times or distances.
Remark 3
The initial route of each CAV is computed based on the prevailing traffic conditions at the time the vehicle starts its trip, to minimize its own travel time. When rerouting is triggered for a group of CAVs, the system planner reoptimizes their routes to minimize their total travel time, thus enhancing overall traffic efficiency. This approach accounts for the diverse departure times of CAVs and ensures that the minimal set of CAVs is rerouted only when necessary, balancing CAVs’ flexibility, transit reliability, and system-wide efficiency.
V Simulation Study
This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed routing approach in a mixed traffic network. We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis.
Figure 4: Visualization of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue. DLs are highlighted in grey, with bus stops marked in yellow.
TABLE I: Departure times of buses
Bus Index 1 2 3 4 5 6 7 8 9 10
Departure Time
[minute] 00 $6$ $12$ $18$ $24$ $30$ $36$ $42$ $48$ $54$
V-A Simulation Setup
As shown in Fig. 4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $5$ bus stops, with the road topology matching that shown in Fig. 1. We consider that buses follow a predefined route along the joint DLs, traveling from the origin stop (node $7$ in Fig. 1) to the destination stop (node $15$). All other edges in the network are designated as GPLs. We assume sensors are installed at each road intersection, allowing for real-time observation of the number of vehicles entering each edge within a predefined monitoring time window. Specifically, the monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
In the road network, we consider $10$ buses operating on the same route from the origin stop to the destination stop, with departures scheduled every $6$ minutes, as given in Table I. The scheduled on-time arrivals at bus stops are based on the free-flow travel time, with an added disturbance margin of $30$ seconds to account for potential delays caused by traffic signals at intersections between bus stops. CAVs and HVs enter the network at rates of $80$ and $120$ vehicles per minute, respectively. The OD pair of each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
Figure 5: The dynamics of the accumulated bus delay.
Figure 6: The dynamics of the accumulated CAV delay.
We compare the proposed coordinated routing approach with two baseline methods: Static Route Planning (SRP) without rerouting and Dynamic Route Planning (DRP) without rerouting. For ease of reference, we refer to the two methods as SRP and DRP without rerouting, respectively. In the SRP without rerouting approach, each CAV is assigned a shortest route based on the static free-flow travel times on the network edges and adheres to this route during its entire trip without any rerouting. In contrast, the DRP without rerouting incorporates real-time traffic data to update travel times on each edge and provides the shortest route based on the congestion conditions when the CAV enters the network. Similar to the SRP approach, no rerouting occurs for CAVs after deployment. The classic Dijkstra algorithm is employed to compute the shortest path for each CAV in the SRP without a rerouting scheme, while a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path in the DRP and proposed rerouting approaches. The weight graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the predicted real-time travel conditions.
V-B Results and Analysis
We track the accumulated bus and CAV delays, measured as the total additional travel time incurred by all buses and CAVs relative to their free-flow travel times along each edge, over a 3500-second simulation horizon. The comparative results are presented in Fig. 5 and Fig. 6, respectively. As shown in Fig. 5, buses experience significantly greater delays under a static shortest-route assignment. In contrast, the proposed method significantly reduces bus delays by integrating real-time congestion estimation with efficient CAV rerouting. The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence. Similar trends are observed for CAV delays, as shown in Fig. 6. Although rerouting occasionally introduces detours for individual vehicles, the dynamic strategy yields lower overall delays. This improvement arises from proactively directing CAVs away from congestion-prone edges, especially those affected by stationary buses, thereby enhancing flow continuity throughout the network.
Figure 7: The accumulated bus delay at each stop.
Figure 8: The accumulated delay of each bus at experienced stops.
Fig. 7 shows the accumulated trip delay for $10$ buses at each stop, comparing the DRP method without rerouting and the proposed approach. Bus stops indexed as $1$, $2$, $3$, and $4$ correspond to nodes $9$, $11$, $13$, and $15$ in Fig. 1. As is shown, the proposed method significantly reduces bus delays at all stops, achieving approximately a $53.6\%$ reduction in delay per stop compared to the DRP method. Fig. 8 further illustrates the accumulated delay experienced by each bus across all stops along the route. It is shown that our rerouting strategy consistently reduces delays for individual buses upon stop arrival, enhancing schedule adherence and overall transit reliability. Additionally, buses $5$ to $8$ experience relatively higher delays compared to buses $1$, $2$, $9$, and $10$, indicating that buses entering the network earlier or later encounter less congestion, aligning with the observed distribution of CAV and HV volumes in the transportation system.
Figure 9: Comparison of vehicle delays with and without using joint DLs.
To demonstrate the advantages of implementing joint DLs, Fig. 9 presents the average trip delays experienced by HVs and CAVs, comparing scenarios with and without the use of joint DLs. In the baseline case, all vehicles utilize GPLs exclusively. As illustrated in the figure, both HVs and CAVs benefit from substantially reduced trip delays under the joint DL configuration, saving around $38\%$ and $40\%$ trip times, respectively, indicating a marked enhancement in overall traffic performance and system-level efficiency.
The results presented above validate the effectiveness of the proposed rerouting strategy in preserving the transit priority of buses and enhancing their schedule adherence. By sharing joint DLs with buses, CAVs also benefit from the dynamic rerouting strategy through reduced trip delays. These outcomes demonstrate that the congestion-aware coordinated routing approach improves overall travel efficiency across the system.
VI Conclusions
In this paper, we proposed a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment. A central element of our approach was a real-time traffic monitoring mechanism that estimated traffic flow and travel time on each road segment using sensor-derived data. Leveraging these estimates, we designed a coordinated scheme that identified a minimal subset of CAVs for rerouting and assigned them optimized paths reflecting current traffic conditions. By dynamically updating CAV routes in response to predicted congestion along dedicated lanes, our method effectively minimized disruptions to transit operations. Simulation experiments conducted in SUMO confirmed that the proposed strategy improved both transit reliability and CAV performance under diverse traffic scenarios. In future work, we plan to extend this framework to accommodate multi-modal transportation systems and assess their scalability in large-scale urban networks.

### v2

@@
 Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2].
-The adoption of CAVs is projected to substantially increase in the coming decades; the transition towards a fully autonomous fleet will occur gradually [3].
-Transportation systems will experience prolonged periods of mixed traffic conditions where CAVs and human-driven vehicles (HVs) coexist.
-Understanding how to optimize the efficiency under increasing penetration rates of CAVs during this transitional period is critical [4, 5, 6].
-Extensive experimental studies have indicated that heterogeneous traffic environments involving CAVs can effectively reduce total travel times [7, 8, 9].
-Implementing suitable right-of-way (ROW) management strategies is essential for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transitional phase.
+While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [3, 4].
+During this period, transportation networks will experience prolonged phases of mixed traffic, where CAVs and human-driven vehicles (HVs) coexist.
+Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [5, 6, 7].
+Extensive research and experimental studies have shown that heterogeneous traffic environments involving CAVs can significantly reduce travel times [8, 9, 10].
+Among various approaches, implementing suitable right-of-way management is pivotal for ensuring the seamless integration of CAVs into existing transportation networks and for fully leveraging their benefits throughout this transition.
 CAV dedicated lane (DL) has been proposed as an effective strategy to enhance the benefits that CAVs can bring to transportation systems [11].
@@
 Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [17].
-While these lanes provide dedicated driving environments for transit operations, the reserved space cannot taken by other non-transit vehicle lower the traffic efficiency  [17].
+While these lanes provide dedicated driving environments for transit operations, the reserved space cannot be taken by other non-transit vehicles, lowering the traffic efficiency [18].
 Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability.
@@
 explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [20].
-A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, were investigated in [11].
+A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, was investigated in [12].
 While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
-In this paper, we propose to develop a dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses.
+In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses.
 The goal is to prioritize the buses’ adherence to their scheduled timeliness while taking advantage of the joint DLs for CAVs to improve road capacity, reduce congestion, and enhance system-wide efficiency.
 By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs.
-The contributions of this work are summarized as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules.
+The main contributions of this work are as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules.
 (ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected.
-Simulation performed in SUMO demonstrates the effectiveness of the proposed approach, indicating significant improvements in traffic efficiency through coordinated use of joint DLs.
+Simulation performed in SUMO based on realistic traffic network data in San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
 II Problem Statement
-Figure 1: The road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
+Figure 1: An example illustrating the road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
 Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges.
-Taking the road network of Manhattan as an example, as shown in Fig.
+As illustrated in Fig.
 1, the node set $\mathcal{V}$ is comprised of two subsets, denoted as $\mathcal{V}\!=\!\mathcal{S}\cup\mathcal{I}$, where $\mathcal{S}$ represents the set of bus stations and $\mathcal{I}$ represents the set of road intersections.
@@
 With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses.
-Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, thus improving road capacity while avoiding disturbance caused by other HVs on GPLs.
+Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, hence improving road capacity while avoiding disturbance caused by other HVs on GPLs.
 Figure 2: Illustration of the general-purpose lane and dedicated lane.
@@
 $P_{i}(t_{\text{sys}})=\big\{(v_{i,1},v_{i,2}),(v_{i,2},v_{i,3}),\dots,(v_{i,n_{i}},d_{i})\big\},$
-where $v_{i,\ell}\!\in\!\mathcal{I}$, $\ell\!=\!1,\dots,n_{i}$, denotes the $\ell$th road intersection along the planned route of CAV $i$ toward its destination.
-To alleviate traffic congestion on DLs that could lead to delays in bus arrivals at their stops, the cumulative flow of CAVs entering each edge on the DLs is actively monitored.
-Specifically, let $\tau_{i,\ell}^{0}$ represent the free-flow travel time on edge $(v_{i,\ell},v_{i,\ell+1})$, and $t_{i,\ell}$ denote the arrival time of CAV $i$ at intersection $v_{i,\ell}$.
-Thus, the arrival time of CAV $i$ at the next intersection $v_{i,\ell+1}$ is of the form
+[…]
 $t_{i,\ell+1}=t_{i,\ell}+\tau_{i,\ell}^{0}.$ (6)
@@
 $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})=\frac{\sum_{i\in\mathcal{N}^{\text{cav}}}\!I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})+\Delta N^{\text{hv}}_{(v,v^{\prime})}(t_{\text{sys}})}{2\Delta{\tilde{T}}}$, (10)
-where the indicated function $I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})$ characterizes the number of CAVs anticipated to traverse the edge $(v,v^{\prime})$ within the monitoring time window.
-Mathematically, it is denoted as
+[…]
 $I_{(v,v^{\prime})}(t_{\text{sys}},\Delta\tilde{T})=\begin{cases}1,&\text{if }t_{i,\ell+1}\!\in\!\mathcal{M}(t_{\text{sys}},\Delta\tilde{T}),\\ &(v_{i,\ell+1},v_{i,\ell+2})\!=\!(v,v^{\prime}),\\ \\ 0,&\text{otherwise},\end{cases}$ (11)
@@
 A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges.
-With different emphasis, the sizes of the two monitoring windows are not necessarily to be identical.
+With a different emphasis, the sizes of the two monitoring windows are not necessarily identical.
 IV Dynamic Rerouting Approach
@@
 Let $\hat{t}_{b,k}$ denote the estimated arrival time of bus $b\!\in\!\mathcal{N}^{\text{b}}$ at its $k$th road intersection $v_{b,k}$.
-Upon arriving at its $(k\!-\!1)$th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
+Upon arriving at its (k1)(k\!-\!1)th road intersection at time $t_{b,k-1}\!=\!t_{\text{sys}}$, the route ahead is denoted as
 $P_{b}(t_{\text{sys}})=\big\{(v_{b,k},v_{b,k+1}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big\},$ (13)
@@
 Figure 3: Illustration of the rerouting scheme at each intersection in DLs.
-To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $[t_{b,k-1},\hat{t}_{b,k}]$.
+To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$.
 The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected.
@@
 Remark 3
+Solving the above route optimization problem requires real-time traffic data collected from existing sensor infrastructures, such as loop detectors and cameras.
+These data are aggregated within monitoring windows to compute vehicles’ actual travel times on the current edges and to estimate their travel times on future edges.
+Incorporating travel time estimations enables long-horizon congestion prediction and route adjustment in advance, thus preventing vehicles from being trapped in severe congestion.
+Remark 4
 The initial route of each CAV is computed based on the prevailing traffic conditions at the time the vehicle starts its trip, to minimize its own travel time.
@@
 V Simulation Study
-This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed routing approach in a mixed traffic network.
+This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed approach in a mixed traffic network.
 We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis.
-Figure 4: Visualization of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue.
+The code implementation is available online.
+[^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency
+Figure 4: Illustration of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue.
 DLs are highlighted in grey, with bus stops marked in yellow.
-TABLE I: Departure times of buses
-Bus Index 1 2 3 4 5 6 7 8 9 10
-Departure Time
-[minute] 00 $6$ $12$ $18$ $24$ $30$ $36$ $42$ $48$ $54$
+TABLE I: Scheduled bus arrival times
+Stop Station 1 Station 2 Station 3 Destination
+Time [mm:ss] 01:03 03:46 07:39 09:27
 V-A Simulation Setup
 As shown in Fig.
-4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $5$ bus stops, with the road topology matching that shown in Fig.
+4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $3$ bus stops, with the road topology matching that shown in Fig.
 1.
-We consider that buses follow a predefined route along the joint DLs, traveling from the origin stop (node $7$ in Fig.
-1) to the destination stop (node $15$).
+This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street, which includes one lane of DL.
+The roadway geometry and signal settings were obtained from OpenStreetMap.
+We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig.
+1) to the destination (node $15$).
 All other edges in the network are designated as GPLs.
-We assume sensors are installed at each road intersection, allowing for real-time observation of the number of vehicles entering each edge within a predefined monitoring time window.
-Specifically, the monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
-In the road network, we consider $10$ buses operating on the same route from the origin stop to the destination stop, with departures scheduled every $6$ minutes, as given in Table I.
-The scheduled on-time arrivals at bus stops are based on the free-flow travel time, with an added disturbance margin of $30$ seconds to account for potential delays caused by traffic signals at intersections between bus stops.
-CAVs and HVs enter the network at rates of $80$ and $120$ vehicles per minute, respectively.
-The OD pair of each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
+We assume sensors are installed at road intersections, enabling real-time observation of the number of vehicles entering each edge within a predefined monitoring time window.
+The monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
+In the road network, we modeled the workday peak-hour bus service with $10$ buses operating along the same route from the origin stop to the destination stop, departing at intervals of $6$ minutes.
+The scheduled on-time arrivals at bus stops were determined based on free-flow travel times, with an additional $60$ seconds of dwell time at each station and a $30$-second disturbance margin to account for potential delays caused by traffic signals between stops.
+The scheduled arrival times for a representative bus are provided in Table I.
+CAVs and HVs enter the network at rates of $8$ and $20$ vehicles per minute, respectively, corresponding to a $30\%$ CAV penetration rate.
+The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
+We compare the proposed coordinated routing strategy against two baseline methods: (i) static route planning (SRP) without rerouting and (ii) dynamic route planning (DRP) with rerouting.
+In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip.
+In the DRP approach, each CAV is assigned the shortest path derived from real-time traffic data at the entry.
+Then, rerouting is applied if a lower-travel-time path is detected within a 60-second monitoring window.
+In the proposed method, CAVs follow the route recommendation obtained by the predictive control approach described in Section IV.
+Different from our approach, congestion anticipation on future road segments is not incorporated in the SRP and DRP.
+For both SRP and DRP, a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path.
+The weighted graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the real-time traffic conditions.
+Regarding the car-following-model in all three scenarios, CAVs use the intelligent driver model with a short headway ($0.6$ s) to capture automation-enabled efficiency.
+The maximum speed is set to $72\,\text{km/h}$, consistent with corridor speed limits.
+HVs and buses use the Krauss model: HVs with moderate headways (about $1.3$ s), variability to reflect human drivers, and a maximum speed of $72\,\text{km/h}$.
+Buses operate with larger headways ($1.6$ s) and a lower maximum speed of $65\,\text{km/h}$.
+This configuration reflects the capacity benefits of CAVs while realistically modeling human drivers and bus operations.
+V-B Results and Analysis
 Figure 5: The dynamics of the accumulated bus delay.
-Figure 6: The dynamics of the accumulated CAV delay.
-We compare the proposed coordinated routing approach with two baseline methods: Static Route Planning (SRP) without rerouting and Dynamic Route Planning (DRP) without rerouting.
-For ease of reference, we refer to the two methods as SRP and DRP without rerouting, respectively.
-In the SRP without rerouting approach, each CAV is assigned a shortest route based on the static free-flow travel times on the network edges and adheres to this route during its entire trip without any rerouting.
-In contrast, the DRP without rerouting incorporates real-time traffic data to update travel times on each edge and provides the shortest route based on the congestion conditions when the CAV enters the network.
-Similar to the SRP approach, no rerouting occurs for CAVs after deployment.
-The classic Dijkstra algorithm is employed to compute the shortest path for each CAV in the SRP without a rerouting scheme, while a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path in the DRP and proposed rerouting approaches.
-The weight graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the predicted real-time travel conditions.
-V-B Results and Analysis
-We track the accumulated bus and CAV delays, measured as the total additional travel time incurred by all buses and CAVs relative to their free-flow travel times along each edge, over a 3500-second simulation horizon.
+TABLE II: Bus on-time performance and delays at each station.
+Station 1 Station 2 Station 3
+On-time percentage [%]
+   SRP using joint DL 10 30 30
+   DRP using joint DL 50 60 60
+   Proposed method 90 90 90
+We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon.
 The comparative results are presented in Fig.
+5.
+As shown, buses experience significant delays under a static shortest-route assignment.
+The DRP approach reduces delays to some extent, but still, buses experience noticeable delays.
+In contrast, the proposed method significantly reduces bus delays nearly back to zero by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting.
+The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence.
+Table II reports the on-time percentage of $10$ buses at each station.
+The on-time percentage is defined as the fraction of buses arriving within $60$ seconds of their scheduled arrival times as given in Table I.
+It shows that the on-time percentages of the SRP, DRP, and our proposed method are approximately $23\%$, $57\%$, and $90\%$, respectively, demonstrating the superior performance of our method.
+Figure 6: The dynamics of the cumulative CAV travel time.
+Figure 7: The dynamics of the cumulative HV travel time.
+Figure 8: The average bus delay at each stop.
+Fig.
 6 and Fig.
-6, respectively.
-As shown in Fig.
-5, buses experience significantly greater delays under a static shortest-route assignment.
-In contrast, the proposed method significantly reduces bus delays by integrating real-time congestion estimation with efficient CAV rerouting.
-The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence.
-Similar trends are observed for CAV delays, as shown in Fig.
-6.
-Although rerouting occasionally introduces detours for individual vehicles, the dynamic strategy yields lower overall delays.
-This improvement arises from proactively directing CAVs away from congestion-prone edges, especially those affected by stationary buses, thereby enhancing flow continuity throughout the network.
-Figure 7: The accumulated bus delay at each stop.
+7 show the total travel time of CAVs and HVs over the simulation horizon.
+The proposed method achieves the lowest overall travel time for both CAVs and HVs.
+This improvement arises from proactively directing CAVs from congestion-prone edges, especially those affected by dwelling buses, thereby enhancing flow continuity throughout the network.
+Note that the DRP approach results in longer travel times for CAVs compared to both the SRP and the proposed method.
+Although for HVs, DRP initially reduces travel times, after $1800$ seconds, the travel times increase and even exceed those under SRP.
+These observations are common in transportation routing problems, as CAVs simultaneously identify and switch to shorter routes, which in turn leads to new and often more severe congestion on those routes.
 Figure 9: The accumulated delay of each bus at experienced stops.
 Fig.
-7 shows the accumulated trip delay for $10$ buses at each stop, comparing the DRP method without rerouting and the proposed approach.
-Bus stops indexed as $1$, $2$, $3$, and $4$ correspond to nodes $9$, $11$, $13$, and $15$ in Fig.
+8 shows the average trip delay for $10$ buses at each station, compared between the SRP method using the joint DL and the proposed approach.
+Bus stations indexed as $1$, 22, and $3$ correspond to nodes $9$, $11$, and $13$ in Fig.
 1.
-As is shown, the proposed method significantly reduces bus delays at all stops, achieving approximately a $53.6\%$ reduction in delay per stop compared to the DRP method.
+As we can observe, the proposed method reduces bus delays significantly at all stations, achieving an average reduction of about $90.7\%$ per station compared with the SRP method.
 Fig.
-8 further illustrates the accumulated delay experienced by each bus across all stops along the route.
-It is shown that our rerouting strategy consistently reduces delays for individual buses upon stop arrival, enhancing schedule adherence and overall transit reliability.
-Additionally, buses $5$ to $8$ experience relatively higher delays compared to buses $1$, $2$, $9$, and $10$, indicating that buses entering the network earlier or later encounter less congestion, aligning with the observed distribution of CAV and HV volumes in the transportation system.
-Figure 9: Comparison of vehicle delays with and without using joint DLs.
-To demonstrate the advantages of implementing joint DLs, Fig.
-9 presents the average trip delays experienced by HVs and CAVs, comparing scenarios with and without the use of joint DLs.
-In the baseline case, all vehicles utilize GPLs exclusively.
-As illustrated in the figure, both HVs and CAVs benefit from substantially reduced trip delays under the joint DL configuration, saving around $38\%$ and $40\%$ trip times, respectively, indicating a marked enhancement in overall traffic performance and system-level efficiency.
-The results presented above validate the effectiveness of the proposed rerouting strategy in preserving the transit priority of buses and enhancing their schedule adherence.
-By sharing joint DLs with buses, CAVs also benefit from the dynamic rerouting strategy through reduced trip delays.
-These outcomes demonstrate that the congestion-aware coordinated routing approach improves overall travel efficiency across the system.
+9 further depicts the accumulated delay experienced by each bus across all stations along the route, showing that our rerouting strategy consistently reduces delays for individual buses at stop arrivals, thereby improving schedule adherence and overall transit reliability.
+Note that bus $6$ experiences relatively higher delays than others, mainly due to the great number of vehicles in the system.
+This result indicates that buses entering the network earlier or later experience less congestion.
+Such a pattern also aligns with the observed distribution of CAVs and HVs volumes in the system.
+Figure 10: The average trip delay under different CAV penetration rates.
+We further evaluate the efficiency of the proposed method under different CAV penetration rates.
+Fig.
+10 presents the average trip delays experienced by buses, HVs, and CAVs under penetration levels of $10\%$, $30\%$, and $50\%$, with the total travel demand set to $2,000$ vehicles per hour.
+The results show that HV delays steadily decrease as CAV penetration increases, while CAV and bus delays exhibit little variation between $30\%$ and $50\%$ penetration.
+These findings show that the benefits of increased penetration saturate beyond a certain threshold, with limited additional travel time savings.
+Figure 11: 90th-percentile travel time comparison with and without joint DL.
+To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without using joint DL.
+In this scenario, we replicate the setting of a real dedicated bus lane by allowing only buses to use the lane, thereby ensuring that buses experience no delay, while CAVs follow their initially assigned shortest routes.
+The results show that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of significantly increasing bus travel times.
+In contrast, the proposed method maintains bus travel times close to those in the dedicated bus lane scenario, while further reducing CAV and HV travel times compared to SRP with joint DL.
+Given the above results, we validate the effectiveness of the proposed rerouting strategy in preserving bus transit priority and improving schedule adherence.
+Through the joint use of DLs with buses, both CAVs and HVs benefit from the dynamic rerouting with reduced trip delays.
+These findings demonstrate that the congestion-aware coordinated routing approach improves travel efficiency across the entire system.
 VI Conclusions
-In this paper, we proposed a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment.
-A central element of our approach was a real-time traffic monitoring mechanism that estimated traffic flow and travel time on each road segment using sensor-derived data.
-Leveraging these estimates, we designed a coordinated scheme that identified a minimal subset of CAVs for rerouting and assigned them optimized paths reflecting current traffic conditions.
-By dynamically updating CAV routes in response to predicted congestion along dedicated lanes, our method effectively minimized disruptions to transit operations.
-Simulation experiments conducted in SUMO confirmed that the proposed strategy improved both transit reliability and CAV performance under diverse traffic scenarios.
-In future work, we plan to extend this framework to accommodate multi-modal transportation systems and assess their scalability in large-scale urban networks.
+In this paper, we presented a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment.
+A central element of our approach was a real-time traffic monitoring system that estimates traffic flow conditions on each road segment using sensor data.
+Leveraging these estimates, we developed a coordination scheme that identifies a minimal subset of CAVs for rerouting and assigns them optimized paths reflecting not only current but also anticipated potential congestion.
+By dynamically updating CAV routes in response to time-varying travel environments along DLs, our method effectively mitigates disruptions to transit operations.
+Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance under diverse traffic scenarios.
+Future work will focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban networks.

### v3

@@
 Abstract
-In this paper, we propose a coordinated routing strategy aimed at improving bus schedule adherence and enhancing travel efficiency for connected and automated vehicles (CAVs) operating within a mixed-traffic urban network.
-Our approach capitalizes on the existence of dedicated lanes for buses and CAVs, leveraging real-time traffic data to dynamically reroute CAVs in anticipation of congestion.
-By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, we enable the system to proactively adjust CAV routes when potential interference with bus operations is detected.
-This coordination mitigates delays affecting transit services and reduces travel time for CAVs.
-We evaluate the proposed strategy through simulation studies conducted in the SUMO.
-The results demonstrate significant improvements in both transit reliability and CAV operational performance across a range of traffic conditions.
+This paper proposes a coordinated routing approach that investigates the use of connected and automated vehicles (CAVs) in dedicated bus lanes.
+The aim is to improve bus schedule adherence while enhancing the travel efficiency of CAVs during the transitional phase of mixed traffic environments.
+Our approach utilizes real-time traffic data to dynamically reroute CAVs in anticipation of congestion.
+By continuously monitoring traffic conditions on dedicated lanes and tracking the real-time positions of buses, the system adjusts CAV routes in advance to avoid potential interference with operating buses.
+This cooperation reduces CAV travel times and minimizes delays that impact transit services.
+The proposed strategy is validated using microscopic traffic simulations in SUMO.
+The results demonstrate significant improvements in both transit on-time performance and CAV travel efficiency across a range of traffic conditions.
 I Introduction
 Connected and automated vehicles (CAVs) have emerged as a central element in the evolution of modern transportation systems.
-Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [1, 2].
+Advances in vehicle-to-vehicle, vehicle-to-infrastructure, and vehicle-to-everything communication technologies highlight the significant potential of CAVs in improving traffic efficiency, reducing energy consumption, and enhancing road safety [12, 7, 8].
 While the adoption of CAVs is projected to grow substantially over the coming decades, the transition towards a fully autonomous fleet is anticipated to unfold gradually [13, 1].
-During this period, transportation networks will experience prolonged phases of mixed traffic, where CAVs and human-driven vehicles (HVs) coexist.
 Understanding how to enhance the traffic efficiency under increasing penetration rates of CAVs is therefore critical [21, 11, 17].
@@
 By providing a separate driving environment, DLs enable CAVs to operate with shorter reaction times and smaller headways, thus increasing lane capacity by up to three times compared to general-purpose lanes (GPLs) [4].
-In addition, DLs isolate CAVs from the mixed traffic network, offering a safer driving environment for both CAVs and HVs.
-This contributes to reducing vehicles’ lane changing and overtaking maneuvers, which are among the leading causes of collisions and delays in mixed traffic environments [13].
-Despite the promising benefits of DLs, the high costs and limited feasibility of constructing new lanes in urban areas, due to substantial construction expenses and restricted urban space, often outweigh the advantages of CAV DLs [14].
+However, in urban settings, the substantial construction costs and limited spatial availability often constrain the feasibility of implementing new CAV DLs [14].
 In this context, converting an existing GPL into a CAV DL tends to be a more practical approach.
 However, it may lead to substantial delays for other types of vehicles using GPLs due to the reduced road capacity [10].
-Recent research has shown that the overall system benefit from converting a GPL to a CAV DL may be lower than maintaining the original configuration [16].
-To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has been proposed: the CAV and bus joint DL.
-Unlike conventional approaches that require constructing new lanes or displacing HVs, joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [17].
-While these lanes provide dedicated driving environments for transit operations, the reserved space cannot be taken by other non-transit vehicles, lowering the traffic efficiency [18].
-Allowing CAVs to access these lanes can enhance lane utilization while maintaining transit reliability.
-In the existing literature, Luo et al.
-developed a framework for the optimal placement of CAV DLs to maximize benefits for all vehicles in the system [19].
-Chen et al.
-explored a space-time modeling framework that demonstrates the mixed-use of DLs can enhance CAV travel efficiency and alleviate the overall traffic congestion [20].
-A systematic framework investigating a variety of lane policies, such as exclusive automated vehicle lanes, regular vehicle lanes, and mixed-use lanes, was investigated in [12].
-While the developed approaches provide new insights into the potential benefits of jointly using DLs, these methods remain limited in their ability to dynamically manage CAVs’ routing, particularly in coordinating vehicle behaviors across the network to balance traffic congestion reduction with maintaining reliable bus travel times, especially under fluctuating demand and a dynamic travel environment.
+To address the tradeoff inherent in implementing DLs for CAVs, a promising strategy has emerged: the CAV and bus joint DL.
+Unlike conventional approaches that require constructing new lanes or displacing human-driven vehicles (HVs), joint use of DLs by buses and CAVs repurposes existing DLs, which were originally designed for buses to ensure their on-time arrivals and maintain public transit efficiency [18].
+In the existing literature, Chen et al.
+proposed a space-time modeling framework demonstrating that the mixed use of DLs can enhance CAV travel efficiency and alleviate overall traffic congestion [5].
+More recently, schedule-based joint optimization approaches have been presented to coordinate CAV speeds and bus timetables in Bus Rapid Transit corridors [20].
+In addition, dynamic management strategies have been developed to allocate right-of-way between buses and CAVs through lane-level trajectory planning [19].
+While these studies provide important insights into joint DL operations, they primarily focus on lane-level coordination or schedule-based optimization, and remain limited in dynamically regulating CAV routing at the network level, particularly in balancing congestion mitigation with reliable bus operations under time-varying demand.
 In this paper, we develop an efficient dynamic routing approach to coordinate CAVs traveling in a mixed traffic network, allowing for the DL sharing with buses.
@@
 By leveraging real-time traffic data, the proposed model enables estimation of potential tense flow on network edges and dynamically navigates the rerouting of CAVs to mitigate congestion on DLs.
-The main contributions of this work are as follows: (i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules.
-(ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, enabling real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected.
-Simulation performed in SUMO based on realistic traffic network data in San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
+In summary, the main contributions of this work are twofold.
+(i) We model the rerouting problem of CAVs in a mixed traffic network where buses and CAVs jointly use DLs to fully take advantage of DLs while enhancing buses’ adherence to schedules.
+(ii) We incorporate a dynamic route condition estimation scheme in the rerouting framework, which enables real-time monitoring of traffic flow and timely adjustment of CAV routes when potential DL congestion is detected.
+Simulation performed in the Simulation of Urban MObility (SUMO) environment based on realistic traffic network data of San Francisco demonstrates the effectiveness of the proposed approach, indicating feasibility and significant improvements in traffic efficiency through coordinated use of joint DLs among buses and CAVs.
 II Problem Statement
-Figure 1: An example illustrating the road transportation network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and the dedicated lanes are shown by red edges.
+Figure 1: An example illustrating the road network, where bus stations ($\mathcal{S}$) are shown by red nodes, intersections ($\mathcal{I}$) are shown in black, and DLs are shown by red edges.
 Consider an urban road network represented by a directed graph $G\!=\!(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\!\subseteq\!\mathcal{V}\!\times\!\mathcal{V}$ denotes the set of directed edges.
@@
 Consider each edge consisting of two lanes in the same direction, i.e., joint DL and GPL, where the GPL can be used by both CAVs and HVs, while DLs are reserved only for joint use by buses and CAVs.
-Throughout this paper, we consider three types of vehicles traveling in the transportation network, including HVs, CAVs, and buses.
+We consider three types of vehicles, including HVs, CAVs, and buses.
 The sets of HVs, CAVs, and buses within the monitoring time horizon are denoted by $\mathcal{N}^{\text{hv}}$, $\mathcal{N}^{\text{cav}}$, and $\mathcal{N}^{b}$, respectively.
 Accordingly, the set of all vehicles is represented as $\mathcal{N}\!=\!\mathcal{N}^{\text{hv}}\!\cup\!{\mathcal{N}^{\text{cav}}}\!\cup\!{\mathcal{N}^{b}}$.
-As illustrated in Fig.
-2, buses use only DL and aim to arrive at each bus stop on time, following a specific timetable.
+As given in Fig.
+2, buses use only DLs and follow a specific timetable.
 In contrast, HVs can use only GPLs for travel purposes and route planning.
 With higher flexibility, CAVs can switch between joint DLs and GPLs to improve travel efficiency while minimizing interruptions to buses.
-Compared to employing only GPLs, driving on joint DLs enables CAVs to keep a smaller inter-vehicle distance, hence improving road capacity while avoiding disturbance caused by other HVs on GPLs.
-Figure 2: Illustration of the general-purpose lane and dedicated lane.
-For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented as a sequence of edges as follows:
+For any bus $b\in\mathcal{N}^{\text{b}}$, let $P_{b}$ denote its fixed route, which is represented by a sequence of edges as follows:
 $P_{b}=\big\{(v_{b,1},v_{b,2}),(v_{b,2},v_{b,3}),\dots,(v_{b,N_{b}-1},v_{b,N_{b}})\big\},$ (1)
@@
 Based on real-time traffic conditions, CAV $i$ is assigned an optimal route from its location to the destination.
-The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted by
+The route of CAV $i$ from $o_{i}$ to $d_{i}$ is denoted as
 $P_{o_{i},d_{i}}=\big\{(o_{i},v_{k}),(v_{k},v_{k+1}),\dots,(v_{n},d_{i})\big\},$ (2)
 where $v_{k},v_{k+1},\dots,v_{n}\!\in\!\mathcal{I}$ are road intersections in the pre-planned route of the CAV.
-Note that, the route of each CAV is dynamically updated upon its arrival at each intersection in its route.
+Note that the route of each CAV is dynamically updated upon its arrival at each intersection in its route.
 While CAV routes may traverse both intersections and bus stations, re-optimization of the route toward the destination is triggered only when approaching an intersection.
 HVs are treated as passive participants in the network, following certain routes during their trips.
+Figure 2: A figure illustrating vehicle types, DLs, and GPLs.
 In this paper, we propose to develop a real-time coordinated routing approach for CAVs that fully leverages dedicated lanes, improves bus schedule adherence, and enhances CAV travel efficiency without creating new obstacles for timely bus arrivals at stops.
@@
 III Modeling of Traffic Flow
-This section presents the modeling of road traffic flow in a dynamic transportation network.
+This section presents the modeling of the road traffic flow in a dynamic transportation network.
 We start by introducing the traffic flow on joint DLs to capture the potential delays caused by CAVs.
@@
 The travel time for a vehicle to traverse an edge depends on the real-time traffic flow along the edge.
-As the number of vehicles traveling on an edge increases, traffic flow becomes heavier and travel times tend to rise due to congestion.
-This nonlinear relationship is well captured by the classic Bureau of Public Roads (BPR) function [21].
-Let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$.
-According to the BPR model, the actual travel time to traverse $e$ is denoted as
+Following the classic Bureau of Public Roads (BPR) function [6], let $\tau_{e}^{0}$ denote the free-flow travel time to traverse edge $e$.
+The actual travel time on the edge is then modeled as
 $\tau_{e}=\tau_{e}^{0}\left(1+\alpha\left(\frac{f_{e}}{c_{e}}\right)^{\beta}\right),$ (3)
-where $f_{e}$ represents the real-time flow on edge $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion.
+where $f_{e}$ denotes the real-time flow on $e$, and $c_{e}$ denotes the effective capacity of the same edge, i.e., the maximum flow that can be sustained without causing congestion.
 Both $f_{e}$ and $c_{e}$ are measured in the number of vehicles per unit time.
-In this paper, we assume that all vehicles, regardless of type, travel at the same speed under free-flow conditions.
+We assume that all vehicles, regardless of their types, travel at the same speed under free-flow conditions.
 Additionally, $\alpha$ and $\beta$ are congestion sensitivity parameters.
-Since CAVs react faster and drive closer with shorter headways, they usually smooth traffic flow compared to a fully human-driven lane [19].
-This results in a lower value of $\alpha$.
-In contrast, buses are heavier and less flexible, causing greater congestion when traffic is dense [22].
-Based on these considerations, we adopt the following parameter settings in this study
-$\begin{cases}\alpha\!=\!0.1,\ \beta\!=\!3,&\text{for GPL},\\ \alpha\!=\!0.2,\ \beta\!=\!5,&\text{for joint DL}.\end{cases}$ (4)
 Sensors deployed in the road network enable real-time monitoring and analysis of traffic flow patterns [2].
-In the road network considered, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge.
-For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, the traffic flow monitor window is defined as
+In the road network, we assume that the traffic flow on each edge is measured by sensors located at the starting point of the edge.
+For each edge on the DLs of bus $b\!\in\!\mathcal{N}^{\text{b}}$, i.e., $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}}$, we define the traffic flow monitor window
 $\mathcal{M}(t_{\text{sys}},\Delta T)=\big[t_{\text{sys}}\!-\!\Delta T,\;t_{\text{sys}}\!+\!\Delta T\big],$ (4)
-where $t_{\text{sys}}$ denotes the current system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window where $\mathbb{N}_{+}$ denotes the set of positive integers.
-For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the transportation system at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its routing planning, which is represented as
+where $t_{\text{sys}}$ denotes the system time and $\Delta{T}\in{\mathbb{N}}_{+}$ is a given constant representing the symmetric monitoring window, with $\mathbb{N}_{+}$ denoting the set of positive integers.
+For each CAV $i\!\in\!\mathcal{N}^{\text{cav}}$ in the road network at time $t_{\text{sys}}$, let $P_{i}(t_{\text{sys}})$ denote its route planning, which is denoted as
 $P_{i}(t_{\text{sys}})=\big\{(v_{i,1},v_{i,2}),(v_{i,2},v_{i,3}),\dots,(v_{i,n_{i}},d_{i})\big\},$
@@
 III-B Traffic Flow on General-Purpose Lane
-Observing traffic flow on GPLs is essential to dynamically optimize CAV route planning based on real-time traffic data, thereby enhancing overall traffic efficiency.
 Similarly, the time window for traffic flow monitoring at the starting point of each edge on the GPLs is denoted as
@@
 where $(v,v^{\prime})\!\in\!\mathcal{E}\!\setminus\!\cup_{b\in\mathcal{N}^{\text{cav}}}P_{b}$ denotes an edge located on the GPLs.
-Note that $\Delta{N}^{\text{hv}}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
-Subsequently, according to the BPR model given in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
-$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right).$ (12)
-Here, $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on edge $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow on the edge at time $t_{\text{sys}}$, computed according to (10).
-The congestion sensitivity parameters $\alpha$ and $\beta$ are determined by (4).
+Note that $\Delta{N}^{\text{hv}}_{(v,v^{\prime})}(t_{\text{sys}})$ represents the increase in the number of HVs, whose route choices are not controlled by the system planner but can be monitored.
+Subsequently, by the BPR model in (3), the travel time for CAVs to traverse each edge on GPLs can be denoted by
+$\hat{\tau}_{(v,v^{\prime})}(t_{\text{sys}})=\tau_{(v,v^{\prime})}^{0}\left(1+\alpha\left(\frac{\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})}{c_{(v,v^{\prime})}}\right)^{\beta}\right),$ (11)
+where $\tau_{(v,v^{\prime})}^{0}$ denotes the free-flow travel time on $(v,v^{\prime})$, and $\hat{f}_{(v,v^{\prime})}(t_{\text{sys}})$ is the anticipated traffic flow computed by (9).
 Remark 1
@@
 A larger monitoring window puts greater emphasis on the macroscopic characteristics of traffic flow on network edges.
-With a different emphasis, the sizes of the two monitoring windows are not necessarily identical.
 IV Dynamic Rerouting Approach
@@
 The objective is to ensure that buses adhere to their scheduled arrival times at each stop while enhancing the overall traffic efficiency of CAVs.
-The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and then determining the optimal alternative routes for those CAVs.
+The proposed optimization scheme consists of identifying the set of CAVs to be rerouted and determining the optimal alternative routes for those CAVs.
 IV-A Identification of CAVs for Rerouting
@@
 $\hat{t}_{b,k}=t_{b,k-1}+\tau_{b,k-1}^{0},$ (13)
-where $\tau_{b,k-1}^{0}$ denotes the free-flow travel time from intersection $v_{b,k-1}$ to intersection $v_{b,k}$.
+where $\tau_{b,k-1}^{0}$ is the free-flow travel time from $v_{b,k-1}$ to $v_{b,k}$.
 Recall that the traffic flow $\hat{f}_{b,k}(t_{\text{sys}})$ on edge $(v_{b,k},v_{b,k+1})$ anticipated at time $t_{\text{sys}}$; cf.
@@
 Employing the BPR function in (3), the anticipated travel time on the same edge is as the following form:
-$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right),$ (15)
-where $c_{b,k}$ is the effective capacity of edge $(v_{b,k},v_{b,k+1})$, defined as the maximum traffic flow that the road segment can accommodate under ideal conditions before congestion affects travel times.
-For edges on DLs, the parameters $\alpha$ and $\beta$ are chosen following the second case in (4).
-Figure 3: Illustration of the rerouting scheme at each intersection in DLs.
-To mitigate traffic congestion on each edge $(v_{b,k},v_{b,k+1})\!\in\!{P_{b}(t_{\text{sys}})}$ caused by CAVs, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$.
-The traffic flow on edge $(v_{b,k},v_{b,k+1})$ is monitored starting from time $t_{b,k-1}$, and CAVs are dynamically rerouted to avoid their entry onto the edge $(v_{b,k},v_{b,k+1})$ whenever congestion is detected.
-We make use of Fig.
-3 for better illustration.
+$\hat{\tau}_{b,k}(t_{\text{sys}})=\tau_{b,k}^{0}\left(1+\alpha\left(\frac{\hat{f}_{b,k}(t_{\text{sys}})}{c_{b,k}}\right)^{\beta}\right).$ (14)
+Figure 3: Rerouting scheme at each intersection in DLs.
+As illustrated in Fig.
+3, we define the control horizon for bus $b$ as the time interval $t_{b,k-1},\hat{t}_{b,k}]$.
 For any system time $t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]$, the set of CAVs on edge $(v_{b,k-1},v_{b,k})$ that need to be rerouted, denoted as $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, is expressed as
 $\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})=\big\{i\!\in\!{\mathcal{N}^{\text{cav}}}:I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1\big\}.$ (15)
-As defined (7), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of the intersection $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered.
+As defined (6), $I_{i}(t_{\text{sys}}^{\prime},\Delta{T},v_{b,k})\!=\!1$ indicates that CAV $i$ is within the monitor window of $v_{b,k}$ at time $t_{\text{sys}}^{\prime}$, which represents the time when congestion on edge $(v_{b,k},v_{b,k+1})$ is detected and CAV rerouting is triggered.
 Namely,
-$t_{\text{sys}}^{\prime}\in\big\{t_{\text{sys}}\!\in\![t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!{\tau_{b,k}^{0}}\big\}.$ (17)
-Remark 2
-In the developed rerouting approach, the traffic flow on edge $(v_{b,k},v_{b,k+1})$ is dynamically monitored, and the CAVs on the preceding edge, i.e., $(v_{b,k-1},v_{b,k})$, are controlled before bus $b$ entering $(v_{b,k},v_{b,k+1})$.
-This procedure enables effective congestion mitigation on each edge along the DLs.
-Thus, it is reasonable to employ the free-flow travel time in (14) to estimate the bus’s arrival time at $v_{b,k}$.
+$t_{\text{sys}}^{\prime}\in\left\{t\in[t_{b,k-1},\hat{t}_{b,k}]:\hat{\tau}_{b,k}(t)\geq(1\!+\!\lambda)\tau_{b,k}^{0}\right\}$ (16)
+where $\lambda\!>\!0$ is a small tolerance.
 IV-B Route Optimization for Rerouted CAVs
-To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\geq{\tau_{b,k}^{0}}$ is detected.
-Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to the destination $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$.
-Based on the latest traffic data, including estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
+To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!(1\!+\!\lambda){\tau_{b,k}^{0}}$ is detected.
+Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$.
+Based on the latest traffic data, including the estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
 $\min\ \sum_{i\in\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})}\sum_{(v_{i,\ell},v_{i,\ell+1})\in{P_{i}(t_{\text{sys}}^{\prime}})}\!\!\!\!\!\!\!\hat{\tau}_{i,\ell}(t_{\text{sys}}^{\prime})$ (17)
@@
 Constraints (4)-(7), (13), and (14) provide estimates of traffic flow and travel times for each edge on joint DLs, while constraints (8)-(11) offer travel time estimates for edges on GPLs.
-Moreover, (16) and (17) specify the triggering condition for rerouting and identify the set of CAVs selected for route replanning.
+Moreover, (15) and (16) identify the set of CAVs selected for route replanning and specify the triggering condition for rerouting.
 After all estimated travel times across the network are updated at time $t_{\text{sys}}^{\prime}$, a prediction-aware Dijkstra algorithm [15] is executed to determine the optimal rerouting paths for all identified CAVs, denoted as $P_{i}^{*}(t_{\text{sys}}^{\prime})$, where $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$.
-In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times or distances.
-Remark 3
-Solving the above route optimization problem requires real-time traffic data collected from existing sensor infrastructures, such as loop detectors and cameras.
-These data are aggregated within monitoring windows to compute vehicles’ actual travel times on the current edges and to estimate their travel times on future edges.
-Incorporating travel time estimations enables long-horizon congestion prediction and route adjustment in advance, thus preventing vehicles from being trapped in severe congestion.
+In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times.
 Remark 2
@@
 V Simulation Study
-This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed approach in a mixed traffic network.
-We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis.
+This section presents simulation studies conducted on the SUMO platform to evaluate the effectiveness of the proposed approach.
 The code implementation is available online.
-[^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency
+[^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency.git
+V-A Simulation Setup
 Figure 4: Illustration of the SUMO simulation environment, where the bus is depicted in yellow, CAVs are shown in red, and HVs are shown in blue.
 DLs are highlighted in grey, with bus stops marked in yellow.
-TABLE I: Scheduled bus arrival times
+As shown in Fig.
+4, we construct an urban road network in SUMO with the road topology matching that shown in Fig.
+1.
+This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street.
+The roadway geometry and signal settings were obtained from OpenStreetMap.
+This corridor is representative of typical transit-priority urban arterials with DL operations.
+We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig.
+1) to the destination (node $15$).
+Sensors are assumed at road intersections, with monitoring windows set to $\Delta T\!=\!30$ s for joint DLs and $\Delta\tilde{T}\!=\!60$ s for GPLs.
+TABLE I: Scheduled bus arrival times.
 Stop Station 1 Station 2 Station 3 Destination
 Time [mm:ss] 01:03 03:46 07:39 09:27
-V-A Simulation Setup
-As shown in Fig.
-4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $3$ bus stops, with the road topology matching that shown in Fig.
-1.
-This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street, which includes one lane of DL.
-The roadway geometry and signal settings were obtained from OpenStreetMap.
-We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig.
-1) to the destination (node $15$).
-All other edges in the network are designated as GPLs.
-We assume sensors are installed at road intersections, enabling real-time observation of the number of vehicles entering each edge within a predefined monitoring time window.
-The monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
-In the road network, we modeled the workday peak-hour bus service with $10$ buses operating along the same route from the origin stop to the destination stop, departing at intervals of $6$ minutes.
-The scheduled on-time arrivals at bus stops were determined based on free-flow travel times, with an additional $60$ seconds of dwell time at each station and a $30$-second disturbance margin to account for potential delays caused by traffic signals between stops.
-The scheduled arrival times for a representative bus are provided in Table I.
-CAVs and HVs enter the network at rates of $8$ and $20$ vehicles per minute, respectively, corresponding to a $30\%$ CAV penetration rate.
+We model a peak-hour service with $10$ buses departing every $6$ min from the same origin.
+The timetable in Table I is aligned with the real schedule and generated from free-flow travel times with $60$ s dwell per stop and a $30$ s signal margin.
+CAVs and HVs enter at rates of $8$ and $20$ veh/min, respectively ($30\%$ CAV penetration).
 The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
@@
 In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip.
-In the DRP approach, each CAV is assigned the shortest path derived from real-time traffic data at the entry.
-Then, rerouting is applied if a lower-travel-time path is detected within a 60-second monitoring window.
-In the proposed method, CAVs follow the route recommendation obtained by the predictive control approach described in Section IV.
-Different from our approach, congestion anticipation on future road segments is not incorporated in the SRP and DRP.
-For both SRP and DRP, a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path.
-The weighted graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the real-time traffic conditions.
-Regarding the car-following-model in all three scenarios, CAVs use the intelligent driver model with a short headway ($0.6$ s) to capture automation-enabled efficiency.
-The maximum speed is set to $72\,\text{km/h}$, consistent with corridor speed limits.
-HVs and buses use the Krauss model: HVs with moderate headways (about $1.3$ s), variability to reflect human drivers, and a maximum speed of $72\,\text{km/h}$.
-Buses operate with larger headways ($1.6$ s) and a lower maximum speed of $65\,\text{km/h}$.
-This configuration reflects the capacity benefits of CAVs while realistically modeling human drivers and bus operations.
+In DRP, each CAV is initially assigned the shortest path using real-time traffic data and is rerouted if a lower-travel-time path is detected within a $60$ s monitoring window.
+In the proposed method, CAVs follow the predictive routing strategy described in Section IV.
+For both SRP and DRP, shortest paths are computed using a weighted Dijkstra algorithm, where edge weights are initialized by free-flow travel times and updated dynamically based on real-time traffic conditions.
 V-B Results and Analysis
 Figure 5: The dynamics of the accumulated bus delay.
-TABLE II: Bus on-time performance and delays at each station.
-Station 1 Station 2 Station 3
-On-time percentage [%]
-   SRP using joint DL 10 30 30
-   DRP using joint DL 50 60 60
-   Proposed method 90 90 90
+TABLE II: Comparison of bus on-time percentage.
+Methods Station 1 Station 2 Station 3
+  SRP using joint DL $10\%$ $\% 3030\%$
+  DRP using joint DL $50\%$ $\% 6060\%$
+  Proposed method $\% 9090\%$ $90\%$
 We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon.
@@
 5.
-As shown, buses experience significant delays under a static shortest-route assignment.
-The DRP approach reduces delays to some extent, but still, buses experience noticeable delays.
-In contrast, the proposed method significantly reduces bus delays nearly back to zero by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting.
-The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence.
-Table II reports the on-time percentage of $10$ buses at each station.
-The on-time percentage is defined as the fraction of buses arriving within $60$ seconds of their scheduled arrival times as given in Table I.
-It shows that the on-time percentages of the SRP, DRP, and our proposed method are approximately $23\%$, $57\%$, and $90\%$, respectively, demonstrating the superior performance of our method.
-Figure 6: The dynamics of the cumulative CAV travel time.
-Figure 7: The dynamics of the cumulative HV travel time.
-Figure 8: The average bus delay at each stop.
+As shown, buses experience significant delays under SRP, while DRP provides only partial improvement.
+In contrast, the proposed method nearly eliminates bus delays by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting.
+Table II reports the on-time percentage of the $10$ buses at each station, defined as arrivals within $60$ s of the scheduled times in Table I.
+The SRP, DRP, and proposed method achieve on average $23\%$, $57\%$, and $90\%$ on-time performance, respectively, demonstrating the advantage of the proposed approach.
 Fig.
@@
 7 show the total travel time of CAVs and HVs over the simulation horizon.
-The proposed method achieves the lowest overall travel time for both CAVs and HVs.
-This improvement arises from proactively directing CAVs from congestion-prone edges, especially those affected by dwelling buses, thereby enhancing flow continuity throughout the network.
-Note that the DRP approach results in longer travel times for CAVs compared to both the SRP and the proposed method.
-Although for HVs, DRP initially reduces travel times, after $1800$ seconds, the travel times increase and even exceed those under SRP.
-These observations are common in transportation routing problems, as CAVs simultaneously identify and switch to shorter routes, which in turn leads to new and often more severe congestion on those routes.
-Figure 9: The accumulated delay of each bus at experienced stops.
-Fig.
-8 shows the average trip delay for $10$ buses at each station, compared between the SRP method using the joint DL and the proposed approach.
-Bus stations indexed as $1$, 22, and $3$ correspond to nodes $9$, $11$, and $13$ in Fig.
-1.
-As we can observe, the proposed method reduces bus delays significantly at all stations, achieving an average reduction of about $90.7\%$ per station compared with the SRP method.
-Fig.
-9 further depicts the accumulated delay experienced by each bus across all stations along the route, showing that our rerouting strategy consistently reduces delays for individual buses at stop arrivals, thereby improving schedule adherence and overall transit reliability.
-Note that bus $6$ experiences relatively higher delays than others, mainly due to the great number of vehicles in the system.
-This result indicates that buses entering the network earlier or later experience less congestion.
-Such a pattern also aligns with the observed distribution of CAVs and HVs volumes in the system.
-Figure 10: The average trip delay under different CAV penetration rates.
+The proposed method achieves the lowest overall travel time for both vehicle classes by proactively diverting CAVs from congestion-prone edges, particularly those affected by bus dwell events, thereby improving network flow continuity.
+Note that the DRP approach yields longer CAV travel times than both SRP and the proposed method, a common effect in dynamic routing where vehicles simultaneously shift to shorter paths and induce secondary congestion.
+Figure 6: The dynamics of the cumulative CAV travel time.
+Figure 7: The dynamics of the cumulative HV travel time.
+Figure 8: The trip delay in different CAV penetration.
 We further evaluate the efficiency of the proposed method under different CAV penetration rates.
@@
 These findings show that the benefits of increased penetration saturate beyond a certain threshold, with limited additional travel time savings.
-Figure 11: 90th-percentile travel time comparison with and without joint DL.
-To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without using joint DL.
-In this scenario, we replicate the setting of a real dedicated bus lane by allowing only buses to use the lane, thereby ensuring that buses experience no delay, while CAVs follow their initially assigned shortest routes.
-The results show that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of significantly increasing bus travel times.
-In contrast, the proposed method maintains bus travel times close to those in the dedicated bus lane scenario, while further reducing CAV and HV travel times compared to SRP with joint DL.
+Figure 9: 90th-percentile travel time comparison.
+To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without joint DL access.
+In this setting, we emulate a dedicated bus lane by allowing only buses to use the DL, while CAVs follow their initially assigned shortest routes.
+Fig.
+9 shows the 90th-percentile travel time for each vehicle type under SRP without joint DL, SRP with joint DL, and the proposed method.
+The comparison reveals that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of increased bus travel time.
+In contrast, the proposed method maintains bus travel time close to that of the dedicated-lane scenario while further reducing CAV and HV travel times compared with SRP using joint DL.
 Given the above results, we validate the effectiveness of the proposed rerouting strategy in preserving bus transit priority and improving schedule adherence.
@@
 VI Conclusions
-In this paper, we presented a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment.
-A central element of our approach was a real-time traffic monitoring system that estimates traffic flow conditions on each road segment using sensor data.
-Leveraging these estimates, we developed a coordination scheme that identifies a minimal subset of CAVs for rerouting and assigns them optimized paths reflecting not only current but also anticipated potential congestion.
-By dynamically updating CAV routes in response to time-varying travel environments along DLs, our method effectively mitigates disruptions to transit operations.
-Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance under diverse traffic scenarios.
-Future work will focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban networks.
+In this paper, we proposed a coordinated routing strategy to improve bus schedule adherence while enhancing CAV travel efficiency in mixed-traffic environments.
+The developed approach proactively identifies a minimal set of CAVs for rerouting based on real-time traffic monitoring conditions and anticipated potential congestion.
+Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance across diverse traffic scenarios.
+A potential direction for future research is to focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban road networks.